owl 1/31/2000 Shift Summary


Subject: owl 1/31/2000 Shift Summary
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:45:59 CST


  
             Shift Summary for owl shift, 01/31/00

On Shift: J. LaDue, B. Hsiung

Hours of beam delivered at all intensities:7.0
Hours of beam delivered at nominal intensity:5.0
Hours of KAMI data 5.0

Accounting of lost time: Access to NM2 to insert polyethaline.
Starting and stopping runs.

===============================================================

Runs ended this shift:
15698 - KAMI_HADRON - 38" of Be
15699 - KAMI_HADRON - Ucal activation effect on triggers.
15700 - KAMI_HADRON - Bad Beam structure at the begining
15701 - KAMI_HADRON - 44" Polyethaline.
15702 - KAMI_HADRON - Ucal activation effect on triggers.
15703 - KAMI_HADRON - Ucal activation effect on triggers.

===============================================================

Tapes brought to FCC this shift: none

Accesses during shift: NM2 to put polyethaline

===============================================================

Description of Shift:

run spill time comment
----------------------------------------------------
             01:06 Stop run 15698

             01:30 Doug J, and Jason L go NM2 to put in the
                      polyethaline.

             02:30 Back from access.

             02:40 Run 15700 stopped, Beam structure was bad.

15704 03:46 Start Long Muon run.

15704 06:30 Enable laser trigger.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:46:02 CST